1. A competitive firm faces a price of 136 and a total cost function of TC=7 + q2 + 6q +5
What is this firm's marginal cost function? MC(q)= 14*q+6

2. What quantity should this firm produce? Leave your answer in fraction form (if necessary).
1
Expert's answer
2019-10-22T09:49:31-0400
"T.C=7q^2+6q+5"


"M.C=\frac{d}{dq}(7q^2+6q+5)=14q+6"

Revenue function

"price\times quantity=136q"


"M.R=\frac{d}{dq}(136q)=136"

At optimal quantity,M.C=M.R

"14q+6=136"


"14q=130"


"q=\frac{65}{7}"